A die is tossed twice. A is the event 'the sum of the two numbers uppermost is greater than 5' .B the event 'the sum of thw two numbers uppermost is less than 8'.
Find the probability of these events.

(a ) A

(b ) B

( c) AB

( d) A U B

Hence show that  P(A U B) =P(A) + P(B) - P(A ∩ B) .

Are A and B mutually eclusive?

Grade:

2 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

The correct option is c)

Shilpi Suman
13 years ago

(a) The probability of A taking place, P(A)= 5/9

(b) The probability of B taking place, P(B)= 7/12

(c) P(AB) = 5/18

(d) P(A U B) = 31/36

P( A U B) = 31/36

P(A)= 5/9

P(B)= 7/12

P( A intersection B) = 5/18

As, 5/9 + 7/12 - 5/18= 31/36

Hence, P(A U B) = P(A) + P(B) - P( A intersection B)